OpenStudy (anonymous):

If u can solve this u r genius! (x2 - 1) (x2 + 1)-1 (1 + x4 )-1/2 dx

OpenStudy (anonymous):

We have : (x2 - 1) (x2 + 1) = X^4-(+1)^4 And : -1 (1 + x4 )=-1-x^4 So have two things kill them : -x^4 and x^4 So (x2 - 1) (x2 + 1)-1 (1 + x4 )=-1-x^4 =-2 and in the next -1/2x too . So have : -2-1/2x For answer ! Just this :) Got it ?:) If u have any problem send a massage :)

OpenStudy (anonymous):

It is solved by integration

OpenStudy (anonymous):

You meant integration???

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

it is one of the toughest questions in mathematics !

OpenStudy (anonymous):

Must exactly told us what you meant .....

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

You should have informed us what you exactly meant .... I thought of it a simple algebraic equation....

OpenStudy (anonymous):

but anyways it does not seem to be a algebric equation !

OpenStudy (anonymous):

dx is read to be sign of a differential element applying in integration. anyway, the answer is: